Use elimination to solve: 4x-y=-3
5x+2y=-20

Answer by checkley71(8403)

5X+2Y=-20 NOW MULTIPLY THE OTHERE EQUATION BY 2 & ADD THEM:
8X-2Y=-6
--------------------
13X=-26
X=-26/13
X=-2 ANSWER.
5*-2+2Y=-20
-10+2Y=-20
2Y=-20+10
2Y=-10
Y=-10/2
Y=-5 ANSWER.
PROOF
4*-2-(-5)=-3
-8+5=-3
-3=-3